OpenStudy (loser66):

Consider the trial on which a 3 is first observed in successive rolls of a six-sides die. Let A be the event that 3 is observed of the first trial. Let B be the event that at least 2trials are required to observe a 3. Assuming that each side has probability 1/6, find 1) P(A) 2) P(B) 3) P(A or B) Please, help.

OpenStudy (loser66):

@phi

OpenStudy (phi):

A is prob of rolling a 3 # of successes/ total # = 1/6

OpenStudy (anonymous):

use binomial

OpenStudy (phi):

event B = Pr # of rolls required > 1 which = 1 - Pr(1 roll) = 5/6

OpenStudy (anonymous):

P(x>2) = 1 - P(x<2)

OpenStudy (loser66):

I know the answer, but I don't know how to put it in logic, first of, I have to define A, and B as sets, How? Please

OpenStudy (loser66):

Like A = {x | x...... something I don't know how to write}

OpenStudy (phi):

I have never done this type of problem. if A is a set, the most obvious is { 3 } B, on the other hand is probably infinite.. consisting of e.g.{ 13,23,43,....}

OpenStudy (phi):

but that hardly seems like a useful way to define probabilities.

OpenStudy (loser66):

Oh, I see A= {3} B={\(3^c3,3^c3^c3,3^c3^c3^c3,....\)} Thank you @phi I got it

OpenStudy (loser66):

so that P(A) =1/6 and P(B) =5/6